OpenStudy (anonymous):

A card is drawn randomly from a standard 52-card deck. Find the probability a) that the card is drawn is a 4 b)The card drawn is a face card c) Card drawn not a face card

OpenStudy (popitree):

there are 4 cards with value 4, so probability is 4/52

OpenStudy (popitree):

=1/13, like that procedd with others

OpenStudy (anonymous):

so then would b be 13/52 and c be 39/52 @popitree

OpenStudy (popitree):

how many face cards are there ? face cards are jacks, queens, and kings

OpenStudy (popitree):

13/52 is not correct

OpenStudy (anonymous):

not exactly sure how many face cards there are in a deck it says theres 13 ranks of each of the 4 french groups. Is c correct?

OpenStudy (popitree):

face cards are jack,queen and king, now there are 4 different types diamond, heats, spade and clubs. so each has their own face cards, so it is 4*3 = 12 face cards

OpenStudy (anonymous):

so 3/13? @popitree

OpenStudy (popitree):

12/52 = 3/13 correct